OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

jimthompson5910 (jim_thompson5910):

Part 1: First let's find the slope m = (y2-y1)/(x2-x1) m = ( -8- 4)/(-3-3) m = (-12)/(-6) m = 2 Now use this along with any point -- say the first one (3,4) -- in y = mx+b to find the equation of the line y = mx+ b y = 2x + b ... Plug in m = 2 4 = 2(3) + b ... Plug in x = 3 and y = 4, this is from the first point 4 = 6 + b 4 - 6 = b -2 = b b = -2 So because m = 2 and b = -2 y = mx+b turns into y = 2x-2

OpenStudy (anonymous):

Thank you, I appreciate it.

jimthompson5910 (jim_thompson5910):

you're welcome